OpenStudy (anonymous):

Divide x ^1/2 vy x ^1/3

OpenStudy (anonymous):

vy is supposed to be by

OpenStudy (anonymous):

Use the property that \[{a^k \over a^l} = a^{k-l}\]

OpenStudy (anonymous):

So using that in this case, what do you get for: \[\Large{ x^{\frac{1}{2}} \over x^{\frac{1}{3}}} = ?\]

OpenStudy (saifoo.khan):

\[\huge\color\orange {x^{\frac{1}{2}-\frac{1}{3}}}\]

OpenStudy (saifoo.khan):

now take the LCM & solve.

OpenStudy (anonymous):

low contrast hurts.

OpenStudy (saifoo.khan):

\[\huge \color\purple {x^{\frac{1}{2}-\frac{1}{3}}}\]